the outer portion of a 22m by 30m lawn is to replaced by a walkway of uniform width. if the lawn is reduced to 3/4 of its original area, find the walkway width
Answer by richwmiller(17219)

22*30*1/4
11*15=165
495
(22-x)*(30-x)=495
x^2-52 x+660 = 495
x^2-52 x+165 = 0
x = 26-sqrt(511)
x = 3.3947